If the terminal side of ∅ is in the second quadrant and sin ∅ = 2/3. Find the values of the trigonometric functions of ∅.

1
Expert's answer
2022-02-16T08:06:06-0500

Opposite side =2, hypotenuse = 3


Adjacent"=\sqrt{(3^2-2^2}=\sqrt5"

"Cos \theta=-\frac{√5}{3}\>\>tan\theta=-\frac{2}{√5}"

"csc\,\theta=\frac{3}{2}\>\>\>sec\>\theta=-\frac{3}{√5}"


"cot\>\theta=-\frac{√5}{2}"